# Maximal discs of Weil-Petersson class in $\mathbb{A}\mathrm{d}\mathbb{S}^{2,1}$

### v1

@@ 3.2. Mess map
Here $(\mathbb{D},w)$ denotes the origin in $T_{0}(1)$ and $(\mathbb{D},z)$ represents the conformal structure on $\Sigma$ determined by $\mu$. Note that the first inequality follows from Lemma 3.3 of [15] and the second inequality follows from the definition of $T^{*}T_{0}(1)$. Hence, these imply that the restrictions of maps $F_{+},F_{-},z$ to the boundary are quasisymmetric homeomorphisms of $\mathbb{S}^{1}$ representing points in $T_{0}(1)$ respectively. Since the composition of two quasisymmetric homeomorphisms representing elements in $T_{0}(1)$ is a quasisymmetric homeomorphism in $T_{0}(1)$, the restriction of $z_{\pm}=F_{\pm}\circ z$ to the boundary $\mathbb{D}$ is a quasisymmetric homeomorphism in $T_{0}(1)$. Hence, one can conclude that $\mathrm{Mess}$ maps $T^{*}T_{0}(1)$ into $T_{0}(1)\times T_{0}(1)$.
For given quasisymmetric homeomorphisms $f_{\pm}:\mathbb{S}^{1}\to\mathbb{S}^{1}$ representing two points in $T_{0}(1)$, let us take quasi-conformal extension determined by the corresponding the Beltrami differentials and we denote them by $z_{\pm}:(\mathbb{D},w)\to(\mathbb{D},z_{\pm})$. Now, by Proposition 3.4 and Remark 3.8, for the pair $(f_{+},f_{-})$, there exists a minimal Lagrangian diffeomorphism extension $F:(\mathbb{D},z_{-})\to(\mathbb{D},z_{+})$ of $f_{+}\circ f_{-}^{-1}:\mathbb{S}^{1}\to\mathbb{S}^{1}$. Moreover, for $(\mathbb{D},z)$ denoting the conformal structure of the graph of $F=F_{+}\circ F_{-}^{-1}$ induced from $(\mathbb{D},z_{+})\times(\mathbb{D},z_{-})$, there exist harmonic maps $F_{\pm}:(\mathbb{D},z)\to(\mathbb{D},z_{\pm})$ such that $F|_{\partial{\mathbb{D}}}=f_{+}\circ f_{-}^{-1}$. Here note that $(\mathbb{D},z)$ may not be same as the conformal structure of $\Sigma$ induced from $\mathbb{A}\mathrm{d}\mathbb{S}^{2,1}$. By the above constructions and Lemma 3.3 in [15], we have
(3.16) $\int_{\mathbb{D}}|\mu_{F}|^{2}e^{\psi}\,d^{2}z_{-}<\infty.$
By (2.9), the Beltrami differential $\mu_{F}$ has the following expression
where $\mu_{\pm}=\mu_{F_{\pm}}$ so that
Combining these,
(3.19) $\int_{\mathbb{D}}|\mu_{F}|^{2}e^{\psi}\,d^{2}z_{-}=\,\int_{\Sigma}\left(|\mu_{F}|^{2}\circ F_{-}\right)\,e^{\psi\circ F_{-}}\,\left(\left|(F_{-})_{z}\right|^{2}-\left|(F_{-})_{\bar{z}}\right|^{2}\right)\,d^{2}z\\ =\,\int_{\Sigma}\left(|\mu_{F}|^{2}\circ F_{-}\right)\,e^{\psi\circ F_{-}}\,\left|(F_{-})_{z}\right|^{2}\left(1-\left|\mu_{-}\right|^{2}\right)\,d^{2}z\\ =\,\int_{\Sigma}\frac{4|\mu_{-}|^{2}}{(1+|\mu_{-}|^{2})^{2}}e^{\phi}\left(1-\left|\mu_{-}\right|^{2}\right)\,d^{2}z.$
Since $[\mu_{F}]\in T_{0}(1)$, $|\mu_{F}(z)|\to 0$ as $|z|\to 1$, so that $|\mu_{-}(z)|\to 0$ as $z$ approaches to $\partial\Sigma$ by (3.18). Hence, there exists $C>0$ such that
(3.20) $C<\frac{1-|\mu_{-}|^{2}}{(1+|\mu_{-}|^{2})^{2}}<1\qquad\text{over}\quad\Sigma.$
Hence, by (3.19) and (3.20),
(3.21) $\int_{\mathbb{D}}|\mu_{F}|^{2}e^{\psi}\,d^{2}z<\infty\qquad\text{ if and only if}\qquad\int_{\Sigma}{|\mu_{\pm}|^{2}}e^{\psi}\,d^{2}z\leq\int_{\Sigma}{|\mu_{\pm}|^{2}}e^{\phi}\,d^{2}z<\infty.$
@@
This means that $F_{\pm}|_{\partial{\mathbb{D}}}$ represents points in $T_{0}(1)$. For the given map $f_{\pm}$, which is the same as $z_{\pm}|_{\partial{\mathbb{D}}}$, representing points in $T_{0}(1)$ respectively, we have that $z|_{\partial{\mathbb{D}}}=(F_{\pm}^{-1}\circ z_{\pm})|_{\partial{\mathbb{D}}}$ represent a point in $T_{0}(1)$. Hence, the pair $(\mu_{z},\Phi)$ represents a point in $T^{*}T_{0}(1)$ where $\mu_{z}$ is the Beltrami differential of $z$ and ±$\pm\Phi$ denotes the Hopf differential of $F_{\pm}$. Here the fact $\Phi\in T^{*}T_{0}(1)$ follows from (3.21). In conclusion, the map $\mathrm{Mess}$ maps $(\mu_{z},\Phi)\in T^{*}T_{0}(1)$ to the given $(z_{+},z_{-})$. Hence, the map $\mathrm{Mess}:T^{*}T_{0}(1)\to T_{0}(1)\times T_{0}(1)$ is onto.
To prove the injectivity of $\mathrm{Mess}:T^{*}T_{0}(1)\to T_{0}(1)\times T_{0}(1)$, we consider the following equalities
where $\mu_{+}=\mu_{F_{+}}$. If the map $\mathrm{Mess}$ is not injective, there are two pairs $(a,b)$ and $(a^{\prime},b^{\prime})$ for $(\mu_{z},z^{*}(\mu_{F}))$ satisfying the above equalities for given $\mu_{z_{\pm}}$. Assuming this, we can derive the following equalities
Combing these, we have
(3.24) $a-a^{\prime}=(\bar{a}-\bar{a^{\prime}})bb^{\prime}.$
Hence, if $a\neq a^{\prime}$, we have $|bb^{\prime}|=1$. But, this contradicts to $|bb^{\prime}|<1$. Hence, we must have $a=a^{\prime}$. This also implies that $b=b^{\prime}$ by (3.23). This completes the proof of the injectivity of $\mathrm{Mess}:T^{*}T_{0}(1)\to T_{0}(1)\times T_{0}(1)$.
@@
To complete the proof of the claim, it is enough to show that the differential of $\mathrm{Mess}:T^{*}T_{0}(1)\to T_{0}(1)\times T_{0}(1)$ is an isomorphism by the inverse function theorem. This will be proved in Proposition 4.8. ∎
As in the proof of Theorem 3.10, one can also prove
@@
For any $(\mu,\Phi)\in T^{*}T_{0}(1)$, the anti-holomorphic energy of the associated induced harmonic maps $F_{\pm}$ defines a finite valued functional over $T^{*}T_{0}(1)$.
@@ 4. Symplectic structure on $T^{*}T_{0}(1)$
In this section, we first derive some variational formulae for the deformations of a maximal disc in the 3-dimensional Anti de Sitter space $\mathbb{A}\mathrm{d}\mathbb{S}^{2,1}$. Using these formulae, we prove the relationship of symplectic forms in terms of the map $\mathrm{Mess}$.
4.1. Variational formulae
For a maximal disc $\Sigma\subset\mathbb{A}\mathrm{d}\mathbb{S}^{2,1}$, we denote by $\Sigma^{\epsilon}\subset\mathbb{A}\mathrm{d}\mathbb{S}^{2,1}$ its deformation family for a small real parameter $\epsilon$. In general, such a deformation consists of two parts: one is a deformation of conformal structures on the maximal disc and the deformation of the way of embedding of the maximal disc in $\mathbb{A}\mathrm{d}\mathbb{S}^{2,1}$. Here we allow the conformal structure of the maximal disc, which needs not to be the same as the induced conformal structure determined by the embedding. For a deformation of conformal structure of the maximal disc, we denote it by $\Sigma^{\epsilon}\cong(\mathbb{D},z_{\epsilon})$. The other part of a deformation of $\Sigma^{\epsilon}$, which is determined by the way of embedding, can be described in terms of a deformation of the induced Gauss map $(F_{+},F_{-}):\Sigma\to(\mathbb{D},z_{+})\times(\mathbb{D},z_{-})$.
To deal with such a general situation of a deformation of a maximal disc $\Sigma\subset\mathbb{A}\mathrm{d}\mathbb{S}^{2,1}$, we consider the following diagram
@@ 4.1. Variational formulae
Here $f^{\epsilon}$and $h^{\epsilon}$ denote quasiconformal maps with the corresponding harmonic Beltrami differential $\epsilon\nu_{f}$ and $\epsilon\nu_{h}$ respectively. Hence, the quasi-conformal maps $f^{\epsilon}$, $h^{\epsilon}$ satisfy the following Beltrami differential equations
(4.2) $f^{\epsilon}_{\bar{z}}=\epsilon\nu_{f}f^{\epsilon}_{z},\qquad h^{\epsilon}_{\bar{u}}=\epsilon\nu_{h}h^{\epsilon}_{u}.$
Taking derivative at $\epsilon=0$, we have
(4.3) $\dot{f}_{\bar{z}}=\nu_{f},\qquad\dot{h}_{\bar{u}}=\nu_{h}.$
@@
For $\nu_{h}=\dot{h}_{\bar{z}}$ satisfying $h^{\epsilon}\circ F=H^{\epsilon}$,
@@
From (2.9), we derive the following expression for the Beltrami differential $\mu_{H^{\epsilon}}$ of $H^{\epsilon}$
(4.5) $\mu_{H^{\epsilon}}=\frac{\mu_{F}+F^{*}(\epsilon\nu_{h})}{1+\bar{\mu}_{F}F^{*}(\epsilon\nu_{h})}.$
Taking derivative of the equality (4.5) at $\epsilon=0$, we have $\dot{\mu}_{H}=(1-|\mu_{F}|^{2})F^{*}(\nu_{h})$. From this, the equality (4.4) follows easily. ∎
For a family of quasiconformal maps $f^{\epsilon}$ with $f^{\epsilon}_{\bar{z}}=\epsilon\nu f^{\epsilon}_{z}$ and a smooth family of tensors $\omega^{\epsilon}$ of type $(\ell,m)$, set
@@
If the diagram (4.1) is commutative, that is, $F^{\epsilon}\circ f^{\epsilon}=H^{\epsilon}=h^{\epsilon}\circ F$,
@@
From the commutative diagram in (4.1), we have
(4.7) $\frac{\mu_{F}+\epsilon F^{*}(\nu_{h})}{1+\epsilon\bar{\mu}_{F}F^{*}(\nu_{h})}=\frac{\epsilon\nu_{f}+(f^{\epsilon})^{*}(\mu_{F}^{\epsilon})}{1+\epsilon\bar{\nu}_{f}(f^{\epsilon})^{*}(\mu_{F}^{\epsilon})}.$
Taking derivative $\epsilon=0$ for (4.7), we have
@@
If the diagram (4.1) is commutative and $\nu_{f}$ is a harmonic Beltrami differential, then
(4.9) $\dot{\mu}_{H}=\nu_{f}-\bar{\nu}_{f}\mu_{F}^{2}+L_{\nu_{f}}\mu_{F}.$
Proof.
The equality (4.9) easily follows from (4.4) and (4.6). ∎
Proposition 4.4.
For a family of Hopf differentials $\Phi^{\epsilon}=e^{\psi^{\epsilon}\circ F^{\epsilon}}F^{\epsilon}_{z}\overline{F}^{\epsilon}_{z}$ on $\Sigma^{\epsilon}$ satisfying the commutative diagram (4.1),
@@
From the definition of Lie derivative, we have
@@
From $h^{\epsilon}\circ F=F^{\epsilon}\circ f^{\epsilon}$, we have
@@
$+e^{\psi\circ F}\Big{(}\dot{h}_{\bar{z}}\circ F\overline{F}_{z}\overline{F}_{{z}}-\dot{\bar{f}}_{z}F_{\bar{z}}\overline{F}_{{z}}+\dot{\bar{h}}_{{z}}\circ F{F}_{{z}}{F}_{{z}}-\dot{\bar{f}}_{{z}}F_{{z}}\overline{F}_{\bar{z}}\Big{)}$
$=$ $\ e^{\psi\circ F}F_{z}\overline{F}_{z}\Big{(}\dot{h}_{\bar{z}}\circ F\overline{F}_{\bar{z}}\overline{F}_{z}(F_{z}\overline{F}_{\bar{z}})^{-1}-\dot{\bar{f}}_{{z}}F_{\bar{z}}F_{z}^{-1}\Big{)}$
@@
For the third equality above, we used the following equality for a variation of the hyperbolic metrics
which follows from the Ahlfors’ lemma given in [1]. This completes the proof. ∎
@@
Expressions of the Lie derivatives in (4.6) and (4.10) are given by both the harmonic Beltrami differentials on $\Sigma$ and the pullback by $F$ of the harmonic Beltrami differentials on the image of $F$. This is due to the deformation of (4.1) involve the deformations on the source and target discs of the harmonic map $F$ simultaneously.
For a family of harmonic diffeomorphism $F_{\pm}^{\epsilon}:\Sigma^{\epsilon}\to\mathbb{D}^{\epsilon}_{\pm}:=(\mathbb{D},(z_{\pm})_{\epsilon})$, let us consider the following commutative diagram
@@
such that $h^{\epsilon}_{\pm}\circ F_{\pm}=F^{\epsilon}_{\pm}\circ f^{\epsilon}$. Moreover, we also assume the corresponding equalities to (4.2) and (4.3) for $f^{\epsilon}$ and $h^{\epsilon}_{\pm}$.
@@
For a family of Hopf differentials $\Phi^{\epsilon}=\Phi^{\epsilon}_{+}$ of $F^{\epsilon}_{+}$ satisfying the commutative diagram (4.14),
@@
where $\nu_{\pm}$ denotes the Beltrami differential $\nu_{h_{\pm}}$ and $\mu_{F}=\mu_{F_{+}}$.
@@
By $\Phi_{\pm}=e^{\phi}\bar{\mu}_{F_{\pm}}$ and Proposition 4.4 applied to $F^{\epsilon}_{\pm}$,
@@
The difference of the equalities (4.17) for $\Phi_{\pm}$ with $\Phi_{+}=-\Phi_{-}$ gives the equality (4.15).
To prove (4.16), we apply Proposition 4.2 to $F^{\epsilon}_{\pm}$ and take the sum of them. Then we have
@@
Combining this and its conjugated equality and noting $\overline{F^{*}_{\pm}(\nu_{\pm})}=F^{*}_{\pm}(\bar{\nu}_{\pm})$ gives the equality (4.16). ∎
@@
In the equalities (4.15) and (4.16), the pullback of the Beltrami differentials $\nu_{\pm}$, $\bar{\nu}_{\pm}$ by $F_{\pm}$ and $\mu_{F_{\pm}}$ have the same tensor type $(-1,1)$. But their dependence are different and they play the different roles in the following sense. The Beltrami differential $\mu_{F_{\pm}}$ depends only on $F_{\pm}:\Sigma\to(\mathbb{D},z_{\pm})$ and play the role of a function over $T_{0}(1)$. On the other hand, the pullbacks of $\nu_{\pm}$, $\bar{\nu}_{\pm}$ by $F_{\pm}$ depend on the deformation and play the role of vectors in the tangent bundle of $T_{0}(1)$.

### v2

@@ 3.2. Mess map
Here $\mathbb{D}_{w}$ denotes the origin in $T_{0}(1)$ and $\mathbb{D}_{z}$ represents the conformal structure determined by $\mu$. The first inequality follows from Lemma 3.3 of [18], and the second inequality follows from the definition of $T^{*}T_{0}(1)$. These conditions imply that the restrictions of maps $F_{+},F_{-}$ and $z$ to the boundary are quasisymmetric homeomorphisms of $\mathbb{S}^{1}$ representing points in $T_{0}(1)$ respectively. Since the composition of two quasisymmetric homeomorphisms representing elements in $T_{0}(1)$ is itself a quasisymmetric homeomorphism in $T_{0}(1)$, the restriction of $z_{\pm}=F_{\pm}\circ z$ to the boundary $\mathbb{D}$ is a quasisymmetric homeomorphism in $T_{0}(1)$. Hence, it follows that $\mathrm{Mess}$ maps $T^{*}T_{0}(1)$ into $T_{0}(1)\times T_{0}(1)$.
For given quasisymmetric homeomorphisms $f_{\pm}:\mathbb{S}^{1}\to\mathbb{S}^{1}$ representing two points in $T_{0}(1)$, let us denote by $z_{\pm}:\mathbb{D}_{w}\to\mathbb{D}_{z_{\pm}}$ the corresponding conformal structures equipped with the hyperbolic structures $I_{\pm}$ respectively. By Proposition 3.9, there exists a pair $(\mu_{z},\Phi)$ representing a point in $T^{*}T(1)$ which is mapped to $(I_{+},I_{-})$ by the $\mathrm{Mess}$ map. We need to show that $(\mu_{z},\Phi)$ represents a point in $T^{*}T_{0}(1)$. To prove this, let us consider the quasisymmetric homeomorphism $f_{+}\circ f_{-}^{-1}:\mathbb{S}^{1}\to\mathbb{S}^{1}$. Then, by Proposition 3.4, there exists a minimal Lagrangian diffeomorphism extension $F:\mathbb{D}_{z_{-}}\to\mathbb{D}_{z_{+}}$ of $f_{+}\circ f_{-}^{-1}$. Moreover, there exist harmonic maps $F_{\pm}:\mathbb{D}_{z}\to\mathbb{D}_{z_{\pm}}$ having ±$\pm\Phi$ as the Hopf differentials, respectively. By the construction, we have $F=F_{+}\circ F_{-}^{-1}$ , and it follows that
(3.32) $\int_{\mathbb{D}}|\mu_{F}|^{2}e^{\psi}\,d^{2}z_{-}<\infty$
by Lemma 3.3 in [18]. From this, as in the proof of Proposition 3.8 we have the equivalence
(3.33) $\int_{\mathbb{D}}|\mu_{F}|^{2}e^{\psi}\,d^{2}z_{-}<\infty\qquad\text{ if and only if}\qquad\int_{\mathbb{D}}{|\mu_{\pm}|^{2}}e^{\psi}\,d^{2}z\leq\int_{\mathbb{D}}{|\mu_{\pm}|^{2}}e^{\phi}\,d^{2}z<\infty.$
@@
This means that $F_{\pm}|_{\partial{\mathbb{D}}}$ represents points in $T_{0}(1)$. For the given map $f_{\pm}=z_{\pm}|_{\partial{\mathbb{D}}}$, which represent points in $T_{0}(1)$ respectively, we have that $z|_{\partial{\mathbb{D}}}=(F_{\pm}^{-1}\circ z_{\pm})|_{\partial{\mathbb{D}}}$ represents a point in $T_{0}(1)$. From the condition $\int_{\mathbb{D}}{|\mu_{\pm}|^{2}}e^{\phi}\,d^{2}z<\infty$ in (3.33), it follows that $\Phi\in T^{*}T_{0}(1)$. Hence, the pair $(\mu_{z},\Phi)$ represents a point in $T^{*}T_{0}(1)$. This concludes that the map $\mathrm{Mess}:T^{*}T_{0}(1)\to T_{0}(1)\times T_{0}(1)$ is surjective.
The injectivity of $\mathrm{Mess}:T^{*}T_{0}(1)\to T_{0}(1)\times T_{0}(1)$ can be proved similarly to the proof of Proposition 3.9.
@@
To complete the proof of the claim, it suffices to show that the differential of the map $\mathrm{Mess}:T^{*}T_{0}(1)\to T_{0}(1)\times T_{0}(1)$ is an isomorphism. Then, by the inverse function theorem, the proof will be completed. This will be established in Proposition 4.7. ∎
As in the proofs of Proposition 3.8 and Theorem 3.10, one can similarly establish the following result.
@@
For the induced harmonic maps $F_{\pm}$ associated to the pairs $(\mu,\Phi)$ representing points in $T^{*}T_{0}(1)$, the anti-holomorphic energy of $F_{\pm}$ defines a finite valued functional $E$ over $T^{*}T_{0}(1)$.
Now we consider the following commutative diagram:
(3.34) ${\large T^{*}T_{0}(1)\ \ }$${\ \ T_{0}(1)\times T_{0}(1)}$${T_{0}(1)}$$\scriptstyle{\mathrm{Mess}}$$\scriptstyle{p_{1}}$$\scriptstyle{p_{2}}$
Here the projection $p_{1}$ is defined as follows: for a given pair $(\mu,\Phi)$ representing a point in $T^{*}T_{0}(1)$, there exist harmonic maps $G_{\pm}:\mathbb{D}_{z}\to\mathbb{D}_{w}$ whose Hopf differentials are ±$\pm\Phi$ respectively, by Theorem 3.2 of [21]. Then, $p_{1}$ maps the pair $(\mu,\Phi)$ to the quasisymmetric homeomorphism $h:=(G_{+}\circ G_{-}^{-1})|_{\partial\mathbb{D}}$, which represents an element in $T_{0}(1)$ by the proof of Theorem 3.10. The second projection $p_{2}$ is defined by
(3.35) $p_{2}(\mu_{+},\mu_{-})=[\mu_{+}\star\mu_{-}^{-1}]$
where the operation $\star$ is defined in terms of (2.10). Then, we observe that the above diagram is commutative, that is, $p_{1}=\mathrm{Mess}\circ p_{2}$. This follows from the constructions of these maps and
$G_{+}\circ G_{-}^{-1}|_{\partial\mathbb{D}}=F_{+}\circ F_{-}^{-1}|_{\partial\mathbb{D}}=z_{+}\circ z_{-}^{-1}|_{\partial\mathbb{D}}.$
Remark 3.12.
The Weil-Petersson universal Teichmüller space $T_{0}(1)$ parametrizes the space of the maximal discs of Weil-Petersson class by Theorem 1.10 of [5]. By the definition of maximal conformal embedding $\sigma:\mathbb{D}_{z}\to\mathbb{A}\mathrm{d}\mathbb{S}^{2,1}$, the cotangent bundle $T^{*}T_{0}(1)$ parametrizes the space of maximal conformal embeddings of Weil-Petersson class. Using (3.14) and (3.35), we observe that $T_{0}(1)$-copy of maximal conformal embeddings corresponds to the same maximal disc. In particular, by Remark 3.6, the inverse images by $p_{i}$ for $i=1,2$ of a totally geodesic disc correspond to the image of the zero section of $T^{*}T_{0}(1)$ and the diagonal set in $T_{0}(1)\times T_{0}(1)$, respectively.
Remark 3.13.
Recall that $T_{0}(1)^{\pm}$ is the inverse image under $\mathrm{Mess}$ of the subsets $T_{0}(1)\times\{0\}$ and $\{0\}\times T_{0}(1)$ in $T_{0}(1)\times T_{0}(1)$. Using (3.14) and (3.35), we observe that these spaces $T_{0}(1)^{\pm}$ also parametrize the space of maximal discs of Weil-Petersson class in $\mathbb{A}\mathrm{d}\mathbb{S}^{2,1}$, respectively.
@@ 4. Symplectic structure on $T^{*}T_{0}(1)$
In this section, we first derive some variational formulas for several quantities associated with a maximal disc in $\mathbb{A}\mathrm{d}\mathbb{S}^{2,1}$, considered along the deformations of a maximal conformal embedding. Using these formulas, we establish a relationship between the canonical form of $T^{*}_{0}T(1)$ and the difference of pullbacks of Weil-Petersson symplectic forms from each factor of $T_{0}(1)\times T_{0}(1)$ via the $\mathrm{Mess}$ map.
4.1. Variational formulas
For a maximal conformal embedding $\sigma:\mathbb{D}_{z}\to\mathbb{A}\mathrm{d}\mathbb{S}^{2,1}$, we denote by
(4.1) $\sigma^{\epsilon}:\mathbb{D}_{z^{\epsilon}}\to\mathbb{A}\mathrm{d}\mathbb{S}^{2,1}$
its deformation family for a small real parameter $\epsilon$. In general, such a deformation consists of two parts: one is a deformation of conformal structures on the domain of $\sigma^{\epsilon}$, and the other is a deformation of the maximal conformal embedding into $\mathbb{A}\mathrm{d}\mathbb{S}^{2,1}$. For the deformation of the conformal structure of the domain of $\sigma^{\epsilon}$, we denote it by $\mathbb{D}_{z^{\epsilon}}$. The other part of the deformation of the maximal conformal embeddings is determined by deformation of the induced Gauss map $(F_{+},F_{-}):\mathbb{D}_{z}\to\mathbb{D}_{z_{+}}\times\mathbb{D}_{z_{-}}$.
To address such a general situation involving the deformation of a maximal conformal embedding $\sigma:\mathbb{D}_{z}\to\mathbb{A}\mathrm{d}\mathbb{S}^{2,1}$, we consider the following diagram:
@@ 4.1. Variational formulas
Here $f^{\epsilon}$and $h^{\epsilon}$ denote quasiconformal maps with the corresponding Beltrami differential $\nu^{\epsilon}_{f}$ and $\nu^{\epsilon}_{h}$, respectively. We may assume that $\nu_{f}^{\epsilon}$ and $\nu_{h}^{\epsilon}$ depend analytically on the real parameter $\epsilon$ such that $\nu^{0}_{f}=0$ and $\nu^{0}_{h}=0$. Therefore, the quasi-conformal maps $f^{\epsilon}$, $h^{\epsilon}$ satisfy the following Beltrami differential equations:
(4.3) $f^{\epsilon}_{\bar{z}}=\nu^{\epsilon}_{f}f^{\epsilon}_{z},\qquad h^{\epsilon}_{\bar{u}}=\nu^{\epsilon}_{h}h^{\epsilon}_{u}.$
It has been known that $f^{\epsilon}$ and $h^{\epsilon}$ depend analytically on $\epsilon$ for every fixed $z$. Taking derivative at $\epsilon=0$, we obtain
(4.4) $\dot{f}_{\bar{z}}=\nu_{f},\qquad\dot{h}_{\bar{u}}=\nu_{h}$
where $\nu_{f}$ and $\nu_{h}$ denote the harmonic Beltrami differential given by the derivative of $\nu_{f}^{\epsilon}$ and $\nu^{\epsilon}_{h}$ at $\epsilon=0$ respectively.
@@
For $\nu_{h}=\dot{h}_{\bar{z}}$ satisfying $h^{\epsilon}\circ F=H^{\epsilon}$, we have
@@
From (2.10), we derive the following expression for the Beltrami differential $\mu_{H^{\epsilon}}$ of $H^{\epsilon}$:
(4.6) $\mu_{H^{\epsilon}}=\frac{\mu_{F}+F^{*}(\nu^{\epsilon}_{h})}{1+\bar{\mu}_{F}F^{*}(\nu^{\epsilon}_{h})}.$
Taking the derivative of this expression with respect to $\epsilon$ at $\epsilon=0$, we obtain
$\dot{\mu}_{H}=(1-|\mu_{F}|^{2})F^{*}(\nu_{h}).$
From this, the equality in(4.5) follows directly. ∎
For a family of quasiconformal maps $f^{\epsilon}$ with $f^{\epsilon}_{\bar{z}}=\nu^{\epsilon}f^{\epsilon}_{z}$ and a smooth family of tensors $\omega^{\epsilon}$ of type $(\ell,m)$, set
@@
If the diagram (4.2) is commutative, that is, $F^{\epsilon}\circ f^{\epsilon}=H^{\epsilon}=h^{\epsilon}\circ F$, then
@@
From the commutative diagram in (4.2), we have the following relation:
(4.8) $\frac{\mu_{F}+F^{*}(\nu^{\epsilon}_{h})}{1+\bar{\mu}_{F}F^{*}(\nu^{\epsilon}_{h})}=\frac{\nu^{\epsilon}_{f}+(f^{\epsilon})^{*}(\mu_{F}^{\epsilon})}{1+\bar{\nu}^{\epsilon}_{f}(f^{\epsilon})^{*}(\mu_{F}^{\epsilon})}.$
Taking the derivative $\epsilon=0$ of this equation, we obtain:
@@
For a family of Hopf differentials $\Phi^{\epsilon}=e^{\psi^{\epsilon}\circ F^{\epsilon}}F^{\epsilon}_{z}\overline{F}^{\epsilon}_{z}$ on $\mathbb{D}_{z^{\epsilon}}$ satisfying the commutative diagram (4.2), the Lie derivative of $\Phi$ along the direction of $\nu_{f}$ is given by
@@
From the definition of the Lie derivative, we have
@@
From the relation $h^{\epsilon}\circ F=F^{\epsilon}\circ f^{\epsilon}$, we obtain
@@
$+e^{\psi\circ F}\Big{(}\dot{h}_{\bar{z}}\circ F\overline{F}_{z}\overline{F}_{{z}}-\dot{\bar{f}}_{z}F_{\bar{z}}\overline{F}_{{z}}+\dot{\bar{h}}_{{z}}\circ F{F}_{{z}}{F}_{{z}}-\dot{\bar{f}}_{{z}}F_{{z}}\overline{F}_{\bar{z}}\Big{)}.$
Using the following identity for variations of hyperbolic metrics,
which follows from the Ahlfors’ lemma given in [1], we obtain
$L_{{\nu_{f}}}\,\Phi=$ $\ e^{\psi\circ F}F_{z}\overline{F}_{z}\Big{(}\dot{h}_{\bar{z}}\circ F\overline{F}_{\bar{z}}\overline{F}_{z}(F_{z}\overline{F}_{\bar{z}})^{-1}-\dot{\bar{f}}_{{z}}F_{\bar{z}}F_{z}^{-1}\Big{)}$
@@
This completes the proof. ∎
@@
The expressions for the Lie derivatives in (4.7) and (4.10) are formulated using both the harmonic Beltrami differentials on the source disc $\mathbb{D}_{z}$ and the pullback by $F$ of the harmonic Beltrami differentials on the target disc. This formulation arises because we consider the general case where deformations occur simultaneously on both the source and target discs of the harmonic map $F$
We now consider a deformation of the induced harmonic maps associated with (4.1), given by a family of harmonic diffeomorphisms
$F_{\pm}^{\epsilon}:\mathbb{D}^{\epsilon}\,\longrightarrow\,\mathbb{D}^{\epsilon}_{\pm},$
where $\mathbb{D}^{\epsilon}:=\mathbb{D}_{z^{\epsilon}}$ and $\mathbb{D}^{\epsilon}_{\pm}:=\mathbb{D}_{z_{\pm}^{\epsilon}}$. Combining both cases, we introduce the following commutative diagram:
@@
such that $h^{\epsilon}_{\pm}\circ F_{\pm}=F^{\epsilon}_{\pm}\circ f^{\epsilon}$. Here we use the notations $\mathbb{D}=\mathbb{D}^{0}$ and $\mathbb{D}_{\pm}=\mathbb{D}^{0}_{\pm}$ for simplicity. Moreover, we assume that $f^{\epsilon}$ and $h^{\epsilon}_{\pm}$ satisfy the corresponding equalities to (4.3) and (4.4).
@@
For a family of Hopf differentials $\Phi^{\epsilon}=\Phi^{\epsilon}_{+}$ of $F^{\epsilon}_{+}$ satisfying the commutative diagram (4.14), we have
@@
where $\nu_{\pm}$ denotes the harmonic Beltrami differential $\nu_{h_{\pm}}$ and $\mu_{F}=\mu_{F_{+}}$.
@@
By $\Phi_{\pm}=e^{\phi}\bar{\mu}_{F_{\pm}}$ and applying Proposition 4.3 to $F^{\epsilon}_{\pm}$, we obtain
@@
Taking the difference of these equalities for $\Phi_{\pm}$ with $\Phi_{+}=-\Phi_{-}$, we derive equation (4.15). To prove (4.16), we apply Proposition 4.2 to $F^{\epsilon}_{\pm}$ and take their sum. This gives
@@
By combining this equation with its conjugated equality and noting that $\overline{F^{*}_{\pm}(\nu_{\pm})}=F^{*}_{\pm}(\bar{\nu}_{\pm})$, we obtain equation (4.16). ∎
@@
In the identities (4.15) and (4.16), the pullbacks of the Beltrami differentials $\nu_{\pm}$, $\bar{\nu}_{\pm}$ by $F_{\pm}$, as well as the Beltrami differentials $\mu_{F_{\pm}}$, all share the same tensor type $(-1,1)$. However, their roles and dependencies differ fundamentally. The Beltrami differential $\mu_{F_{\pm}}$ depends solely on the map $F_{\pm}:\mathbb{D}_{z}\to\mathbb{D}_{z_{\pm}}$, and plays the role of a coordinate over $T_{0}(1)$. In contrast, the pullbacks of $\nu_{\pm}$, $\bar{\nu}_{\pm}$ by $F_{\pm}$ arise from the deformation and represent tangent vectors on $T_{0}(1)$, thereby playing the role of vector fields.